# In-Context Reinforcement Learning for Variable Action Spaces

### v3

@@ Abstract
Recently, it has been shown that transformers pre-trained on diverse datasets with multi-episode contexts can generalize to new reinforcement learning tasks in-context. A key limitation of previously proposed models is their reliance on a predefined action space size and structure. The introduction of a new action space often requires data re-collection and model re-training, which can be costly for some applications. In our work, we show that it is possible to mitigate this issue by proposing the Headless-AD model that, despite being trained only once, is capable of generalizing to discrete action spaces of variable size, semantic content and order. By experimenting with Bernoulli and contextual bandits, as well as a gridworld environment, we show that Headless-AD exhibits significant capability to generalize to action spaces it has never encountered, even outperforming specialized models trained for a specific set of actions on several environment configurations.
Machine Learning, ICML
@@ 1 Introduction
Figure 3: Algorithm Distillation Struggles with Novel Action Spaces: Despite its good results on the train action set, AD’s performance diminishes when the action semantics change, either due to a permutation or substitution. It is important to note that augmenting the training data with permuted action sets does not lead to increased performance, signifying that action set invariance should be enforced from a model design standpoint. Additionally, it is impossible to apply a trained AD model to a larger action set. On the graph, the bars are the success rate values on the gridworld environment (described in Section 4.3) obtained after evaluating each of the action sets visualized in Figure 1, averaged over 5 runs. Altered Semantics aggregate the values from the Permuted Train Actions and Sliced Test Actions sets. Altered Size aggregates the values from Test Actions and All Actions. See Section 4.3 for more information about the construction of the action sets.
@@
Creating models that can adapt to new action spaces is essential for building the foundation of decision-making systems in order to enable large-scale pretraining across various environments and address real-world problems (Jain et al., 2020; Chandak et al., 2020; London & Joachims, 2020; Jain et al., 2021). With this in mind, our research focuses on variable discrete action spaces, with the notion of variability illustrated in the Figure 1. In our study, we reveal the limitations of the Algorithm Distillation (AD) model from prior work, such as its diminished performance upon changes in action semantics as well as architectural constraints when handling varying action space sizes (see Figure 3).
@@
Figure 4: Algorithm Regret under Variable Reward Distributions in Bernoulli Bandit. The graph compares regret for Random, Thompson Sampling, and Headless-AD across distinct reward distributions in the Bernoulli Bandit environment, averaged from five seeds. During training, the high reward was $95\%$ more likely to distribute across the odd arms. During testing, it either switched to the even arms or a uniform distribution. Note that Headless-AD maintains high performance in all configurations, proving its ICL capabilities at generalizing to novel tasks, represented by changes in reward distribution. Data is aggregated from bandit problems with $4-20$ arms, reflecting the training conditions.
@@ 2 Algorithm Distillation Struggles with Novel Action Spaces
Figure 5: Algorithm Regret under Increasing Amount of Arms in Bernoulli Bandit. This series of plots shows the regret of Thompson Sampling, AD, and Headless-AD algorithms over evaluation steps in environments with $20-50$ arms, averaged from five seeds with $100$ bandits each. Although Headless-AD has been trained on bandits with up to $20$ arms, it performs well, matching or outperforming other algorithms in larger arm settings without additional training. Note that AD was retrained from scratch for each task with a different number of arms.
@@ 3 Headless-AD
To mitigate the action space limitations of AD, we propose Headless-AD, a new architecture that improves on AD by omitting the final linear layer and incorporating three key modifications. The Headless-AD architecture and data flow is visualized in Figure 2.
@@
Direct Prediction of Action Embeddings: The model output is modified to yield an action embedding rather than a probability distribution over actions. This alteration makes the model independent of action set size and order, granting it permutation invariance. The InfoNCE Contrastive Loss (Oord et al., 2018), diverging from its usual role in representation learning (Jaiswal et al., 2020), serves as a regression objective to reinforce the similarity between the model prediction and the subsequent action in the data. All other actions are treated as negative samples:
$L=-\mathop{\mathbb{E}}\left[{\log{\frac{f(\hat{a}_{t},a_{t})/\tau}{\sum_{a\in A}{f(\hat{a}_{t},a})/\tau}}}\right],$
where $\tau$ is a temperature parameter. We used dot-product as the similarity function $f$.
Random Action Embedding Strategy: To prepare the model for potential new actions during testing, we transition from static learned action embeddings to dynamic embeddings generated by a mapping function $g:~{}\mathbb{A}~{}\rightarrow~{}\mathbb{R}^{n}$, which produces a unique random encoding for each action in a batch at the start of every training step. The mapping is shared across all batch instances and is consistent along the context sequence, i.e., actions with index $i$ in their respective action sets will all be mapped to the same embedding. The random embeddings are constrained as orthogonal vectors within a unit sphere. The details of the sampling process can be found in Appendix F. This approach makes the model deduce action semantics from the interaction history. The orthogonality also clarifies the interpretation of similarity measures as logits because it isolates the similarity of each action, thus preventing overlap in the probability distribution. Moreover, employing random embeddings enhances data variety, which has been demonstrated to boost in-context learning for RL agents (Kirsch et al., 2023; Lu et al., 2023).
@@
Action Set Prompt: To address the model’s lack of awareness of the action space structure, we prepend the input with a sequence of embeddings for all available actions.
The modified input format is thus represented as:
@@
where $N$ is the action set size and $a^{emb}=g(a)$.
We employ two methods to select actions during inference: (1) nearest neighbor selection, and (2) probabilistic sampling based on the similarity to the predicted action embedding. We treat the specific choice of method as a hyperparameter. An illustrative code snippet with Headless-AD’s training procedure can be found in Appendix I.
@@ 4 Experiments
As Headless-AD extends and improves on AD, we check it in two different aspects. Firstly, it should maintain In-Context Learning abilities and thus generalize well to new tasks. Secondly, it should show high performance on action spaces different from the one seen during training. All of the following environments are designed specifically to check both of the above aspects. In our experiments, we use the TinyLLaMa (Zhang et al., 2024) implementation of the transformer model and AdamW optimizer (Loshchilov & Hutter, 2017). All environment specific hyperparameters are listed in Appendix H.
4.1 Bandit
@@
Motivation: This experiment checks Headless-AD’s abilities on a toy task, where the environment does not have a notion of state and returns binary rewards.
Setup 1: In our first experiment, we examined the model’s robustness to distributional shifts in rewards. We used a Bernoulli bandit where each arm is associated with a mean $\mu$ and the reward after pulling an arm $i$ is generated by $\text{Bernoulli}(\mu)$. The training dataset consists of bandits with $4-20$ arms, i.e., different action set sizes. Additionally, to evaluate the ICL capabilities of models, the reward distribution over the arms differed between train and test (Laskin et al., 2022). Specifically, the training data consisted of bandits where in $95\%$ of cases the odd-numbered arms were assigned a random $\mu\in[0.5,1]$ and even-numbered arms received $\mu\in[0,0.5]$. For other $5\%$ of bandits, the ranges were swapped between odd and even arms. The test distribution also consisted of bandits with $4-20$ arms, but the reward distribution was switched either to even arms or was uniform, i.e., all arms were assigned $\mu\in[0,1]$. Learning histories were generated using Thompson Sampling algorithm for a total of $10,000$ bandit instances with $300$ steps each. The evaluation was performed on $100$ bandits in each reward distribution, included $5$ seeds, and the algorithms were rolled out for $300$ steps.
Results and Discussion 1: As depicted in Figure 4, the Headless-AD model demonstrated strong generalization capabilities by almost reaching the performance results set by the traditional Thompson Sampling algorithm under each test distribution.
Setup 2: In our second experiment, we evaluated the transferability of Headless-AD to new action set sizes. Training distribution remained the same as in the previous experiment. Each evaluation dataset consisted of a fixed amount of $20$, $25$, $30$, $40$ and $50$ arms and a uniform distribution of rewards. AD was trained on fixed-size bandits corresponding to each evaluation dataset. The training and evaluation reward distributions are the same as for Headless-AD.
Results and Discussion 2: Figure 5 showcases that Headless-AD can effectively maintain its performance as the action space grows, without necessitating any retraining. Moreover, Headless-AD even outperforms specially trained AD, especially on larger action sets. Note that Headless-AD’s performance curves resemble TS’s performance curves, signifying that it has indeed learned a policy improvement operator generic enough to apply to unseen action set sizes.
@@ 4.2 Contextual Bandit
Motivation: To validate the sustained performance of Headless-AD, we progress to a more complex Multi-Armed Bandit (MAB) extension that integrates states and real-valued rewards.
Setup: Each time step presents a context with arm features modeled as two-dimensional vectors, and rewards are generated with a standard deviation $\sigma=1$. The data is created using a LinUCB (Li et al., 2010) algorithm trained over $300$ steps, on bandits with $4-20$ arms. We used $100$ contextual bandits, $5$ seeds and $300$ evaluation steps.
Results and Discussion: As shown in Figure 6, Headless-AD’s performance is on par with LinUCB across varied arm counts, while also being capable of adept generalization to new environments. While AD also reaches the performance of LinUCB on lower space sizes, it has problems converging on larger action spaces sizes, in addition to requiring a specific retraining. This serves to highlight the advantages of Headless-AD for use in even more complex environments than toy Bernoulli bandits.
@@ 4.3 Darkroom
Setup: The Darkroom environment, inspired by Chevalier-Boisvert et al. (2018) and Jain et al. (2020), consists of a $N\times N$ grid where the agent needs to reach a specific cell for a reward. Action space consists of $3$-step sequences of $5$ atomic actions: up, down, left, right, noop. As a result, the environment offers $5^{3}$ possible actions. The agent earns a reward of $1$ if the trajectory induced by the action sequence passes through a goal cell, after which the episode finishes. Otherwise, the reward is 00. As an observation, the environment offers only the current coordinates of the agent, so the goal information can only be obtained from the agent’s memory. We divide the goals into disjointed sets used for training and testing in order to evaluate Headless-AD’s in-context learning abilities. Furthermore, the action set is randomly split into train and test sets, each including $50$ and $75$ actions respectively, to create five distinct spaces for assessing various generalization aspects. These action sets are visualized in Figure 1.
@@
Train Actions. Comprising the training split, this set represents the actions encountered during model training.
@@
Test Actions. Comprising the test split, this set assesses Headless-AD’s generalization on novel and larger action sets.
@@
All Actions. Combining both training and testing actions, this set contains $125$ actions and is $2.5$ times larger than the training set. Its aim is to challenge Headless-AD to effectively integrate seen and unseen actions.
@@
Permuted Train Actions. Shuffled training set, meant to test the model’s adaptability to reordered action spaces that are comprised of the same actions.
@@
Sliced Test Actions. Tailored to match the training set size, this set contains a slice of the first $50$ actions from the test set. Its aim is to check the models’ generalization abilities on unseen actions while maintaining the action set size.
@@
In scenarios where the action space exceeds the size of the training set, we analyze the performance of an AD model retrained from scratch. The data generation algorithm is Q-learning, executed over $200$ episodes for each environment. Further details on model hyperparameters are available in Appendix H.
Figure 8: Impact of Action Set Prompt Ablation: (a) The inclusion or exclusion of the prompt, which lists all possible action embeddings, does not substantially affect the number of distinct actions explored by the agents. This plot shows an averaged number of actions values across all action sets. (b) The exclusion of the prompt significantly impacts performance in environments like the Bernoulli Bandit, while the effect is less prominent in the Darkroom environment. This plot shows average performance values across all action sets. Performance on the y-axis represents the normalized regret on the Bernoulli Bandit environment, and the success rate on Darkroom. (c) The exclusion of the prompt has an impact on performance that becomes increasingly significant with the growth of the action space. The trend suggests that, as the number of potential actions rises, the ability to strategically sample actions based on the action set prompt becomes increasingly vital for model efficacy.
@@
Results and Discussion: Figure 7 illustrates the generalization abilities of both AD and Headless-AD. First, note that both models maintain their performance when transitioning to novel tasks, thus fulfilling their purpose as ICL-RL models. However, this environment is mainly designed to challenge the models’ generalization abilities on novel action spaces. Following the scope of the action space novelty we set earlier, we study the models’ ability to address changed action semantics and variable action set sizes.
@@
While AD achieves high performance on the train action set, its limitations become evident when the action space is changed. The first limitation is AD’s action set size constraint. The linear layer at the end of its network fixes the size of the output dimension, something that cannot be modified once the model is trained. Increasing the amount of options, as was done in the Test Actions and All Actions sets, requires reinitializing the output dimension and thus retraining the model, which demands additional time and resources. In contrast, once trained, Headless-AD swiftly adapts to changes in the action set size without losing performance.
@@
The second limitation is AD’s reliance on a stationary action space structure. Due to the classifier nature of AD’s network, it learns to associate each output dimension with a specific action meaning. When action semantics change either due to a permutation of a seen action, as in Permuted Train Actions, or due to a substitution of completely new actions, as in Sliced Test Actions, AD’s performance degrades. We checked whether this problem may be solved by biasing the training distribution to have a structure resembling the one during testing. We trained AD on permuted action sets while leaving the data and the model architecture the same. However, as one can see in Figure 3, this training procedure results only in a slightly decreased performance of AD in the Permuted Train setting. An additional graph depicting the performance on each of the (action set, goal type) pairs can be found in Appendix C. Meanwhile, Permuted Train Actions does not pose a challenge for Headless-AD as, by design, it is invariant to specific action order. Most importantly, Headless-AD maintains its performance even when completely new actions are introduced, despite having never seen them in training. We attribute this feature to our success in making the model infer action meaning from the context.

### v5

@@ Abstract
Recently, it has been shown that transformers pre-trained on diverse datasets with multi-episode contexts can generalize to new reinforcement learning tasks in-context. A key limitation of previously proposed models is their reliance on a predefined action space size and structure. The introduction of a new action space often requires data re-collection and model re-training, which can be costly for some applications. In our work, we show that it is possible to mitigate this issue by proposing the Headless-AD model that, despite being trained only once, is capable of generalizing to discrete action spaces of variable size, semantic content and order. By experimenting with Bernoulli and contextual bandits, as well as a gridworld environment, we show that Headless-AD exhibits significant capability to generalize to action spaces it has never encountered, even outperforming specialized models trained for a specific set of actions on several environment configurations. Implementation is available at: https://github.com/corl-team/headless-ad.
Machine Learning, Reinforcement Learning, In-Context Learning, Variable Action Spaces, Transformers
@@ 1 Introduction
Creating models that can adapt to new action spaces is essential for building the foundation of decision-making systems in order to enable large-scale pretraining across various environments and address real-world problems (Jain et al., 2020; Chandak et al., 2020; London & Joachims, 2020; Jain et al., 2021). With this in mind, our research focuses on variable discrete action spaces, with the notion of variability illustrated in the Figure 1. In our study, we reveal the limitations of the Algorithm Distillation model from prior work, such as its diminished performance upon changes in action semantics as well as architectural constraints when handling varying action space sizes (see Figure 3).
@@
loh
@@ 3 Headless-AD
Figure 3: Algorithm Distillation Struggles with Novel Action Spaces: Despite its good results on the train action set, AD’s performance diminishes when the action semantics change, either due to a permutation or substitution. It is important to note that augmenting the training data with permuted action sets does not lead to increased performance, signifying that action set invariance should be enforced from a model design standpoint. Additionally, it is impossible to apply a trained AD model to a larger action set. On the graph, the bars are the success rate values on the Darkroom environment (described in Section 4.3) obtained after evaluating each of the action sets visualized in Figure 1, averaged over 5 runs. Altered Semantics aggregate the values from the Permuted Train Actions and Sliced Test Actions sets. Altered Size aggregates the values from Test Actions and All Actions. See Section 4.3 for more information about the construction of the action sets.
Figure 4: Algorithm Regret under Variable Reward Distributions in Bernoulli Bandit: The graph compares regret for Random, Thompson Sampling, and Headless-AD across distinct reward distributions in the Bernoulli Bandit environment, averaged from five seeds. During training, the high reward was $95\%$ more likely to distribute across the odd arms. During testing, it either switched to the even arms or a uniform distribution. Note that Headless-AD maintains high performance in all configurations, proving its ICL capabilities at generalizing to novel tasks, represented by changes in reward distribution. Data is aggregated from bandit problems with $4-20$ arms, reflecting the training conditions.
@@
To mitigate the action space limitations of AD, we propose Headless-AD, a new architecture that improves on AD by omitting the final linear layer and incorporating three key modifications. The Headless-AD architecture and data flow are visualized in Figure 2.
Random Action Embeddings: To remove the dependence of the model on the pretrained action embeddings, we employed a dynamic mapping function $g:~{}\mathbb{A}~{}\rightarrow~{}\mathbb{R}^{n}$, which produces a unique random encoding for each action in a batch at the start of every training step. The mapping is shared across all batch instances and is consistent along the context sequence, i.e., actions with index $i$ in their respective action sets will all be mapped to the same embedding. During inference, we generated a single set of embeddings at the beginning and used them throughout the evaluation.
The core intuition behind employing random action embeddings is to eliminate any prior knowledge about the structure of the action space within our model. This approach stems from our observation that using learnable embeddings for actions becomes impractical when encountering new actions not seen during training. A new action would lack a pre-trained embedding, and assigning an arbitrary embedding could introduce an undesirable domain shift, as the model would not recognize it. Moreover, allowing the model to learn new embeddings on-the-fly would necessitate extra gradient steps, diverging from our goal of maintaining a zero-shot learning framework. The usage of random action embeddings ensures that the model does not depend on extracting any information from the embeddings themselves but rather relies on interpreting the context provided by historical interactions with the environment. Moreover, employing random embeddings enhances data variety, which has been demonstrated to boost in-context learning for RL agents (Kirsch et al., 2023; Lu et al., 2023).
Figure 5: Algorithm Regret under Increasing Amount of Arms in Bernoulli Bandit: This series of plots shows the regret of Thompson Sampling, AD, and Headless-AD algorithms over evaluation steps in environments with $20-50$ arms, averaged from five seeds with $100$ bandits each. Although Headless-AD has been trained on bandits with up to $20$ arms, it performs well, matching or outperforming other algorithms in larger arm settings without additional training. Note that AD was retrained from scratch for each task with a different number of arms.
We further refined this strategy by constraining the random embeddings to lie on a unit sphere and ensuring their orthogonality (see Appendix H). The choice of a unit sphere normalizes the scale of the embeddings, while orthogonality allows the model to independently adjust the probability assigned to each action. This condition is crucial for preventing unintended probability mass allocation to multiple actions when the model’s prediction vector aligns with one embedding vector. A similar concept is explored by (Elhage et al., 2022) in the context of feature interference.
@@
Direct Prediction of Action Embeddings: The model output is modified to yield an action embedding $\hat{a}^{emb}_{t}$ rather than a probability distribution over actions. This alteration makes the model independent of action set size and order, granting it permutation invariance. The InfoNCE Contrastive Loss (Oord et al., 2018), diverging from its usual role in representation learning (Jaiswal et al., 2020), serves as a regression objective to reinforce the similarity between the model prediction and the subsequent action in the data. All other actions are treated as negative samples. Thus, the objective is
$L=-\mathop{\mathbb{E}}\left[{\log{\frac{f(\hat{a}^{emb}_{t},a^{emb}_{t})/\tau}{\sum_{a\in A}{f(\hat{a}^{emb}_{t},a^{emb}})/\tau}}}\right],$
where $a^{emb}=g(a)$ and $\tau$ is a temperature parameter. We used dot-product as the similarity function $f$.
@@
Action Set Prompt: To address the model’s lack of awareness of the action space structure caused by the two previous changes, we prepend the input with a sequence of embeddings for all available actions.
The modified input format is thus represented as
@@
where $N$ is the action set size. An illustrative code snippet with Headless-AD’s training procedure can be found in Appendix L.
We suggest two methods to select actions during inference: (1) nearest neighbor selection:
$a=\arg\max_{a\in A}f(\hat{a}^{emb},a^{emb})$
and (2) probabilistic sampling based on the similarity to the predicted action embedding:
$P(a_{i})=\frac{f(\hat{a}^{emb},a^{emb}_{i})}{\sum_{a\in A}{f(\hat{a}^{emb},a^{emb}})}.$
We treat the specific choice of method as a hyperparameter.
@@ 4 Experiments
As Headless-AD extends and improves on AD, we checked it in two different aspects. Firstly, it should maintain In-Context Learning abilities and thus generalize well to new tasks. Secondly, it should show high performance on action spaces different from the one seen during training. All of the following environments are designed specifically to check both of the above aspects. In our experiments, we used the TinyLLaMa (Zhang et al., 2024) implementation of the transformer model and AdamW optimizer (Loshchilov & Hutter, 2017). All environment specific hyperparameters are listed in Appendix J.
4.1 Bernoulli Bandit
@@
Motivation: This experiment checked Headless-AD’s abilities on a toy task, where the environment did not have a notion of state and returned binary rewards.
Setup 1: In our first experiment, we examined the model’s robustness to distributional shifts in rewards. We used a Bernoulli bandit where each arm is associated with a mean $\mu$ and the reward after pulling an arm $i$ is generated by $\text{Bernoulli}(\mu)$. The training dataset consisted of bandits with $4-20$ arms, i.e., different action set sizes. Additionally, to evaluate the ICL capabilities of models, the reward distribution over the arms differed between train and test (Laskin et al., 2022). Specifically, the training data consisted of bandits where in $95\%$ of cases the odd-numbered arms were assigned a random $\mu\in[0.5,1]$ and even-numbered arms received $\mu\in[0,0.5]$. For other $5\%$ of bandits, the ranges were swapped between odd and even arms. The test distribution also consisted of bandits with $4-20$ arms, but the reward distribution was switched either to even arms or was uniform, i.e., all arms were assigned $\mu\in[0,1]$. Learning histories were generated using Thompson Sampling algorithm for a total of $10,000$ bandit instances with $300$ steps each. The evaluation was performed on $100$ bandits in each reward distribution, included $5$ seeds, and the algorithms were rolled out for $300$ steps.
Results and Discussion 1: As depicted in Figure 4, the Headless-AD model demonstrates strong generalization capabilities by almost reaching the performance results set by the traditional Thompson Sampling algorithm under each test distribution.
Setup 2: In our second experiment, we evaluated the transferability of Headless-AD to new action set sizes. Training distribution remained the same as in the previous experiment. Each evaluation dataset consisted of a fixed amount of $20$, $25$, $30$, $40$ and $50$ arms and a uniform distribution of rewards. AD was trained on fixed-size bandits corresponding to each evaluation dataset. The training and evaluation reward distributions were the same as for Headless-AD.
Results and Discussion 2: Figure 5 shows that Headless-AD can effectively maintain its performance as the action space grows, without necessitating any retraining. Moreover, Headless-AD even outperforms a specially trained AD, especially on larger action sets. Note that Headless-AD’s performance curves resemble TS’s performance curves, signifying that it has indeed learned a policy improvement operator generic enough to apply to unseen action set sizes.
@@ 4.2 Contextual Bandit
Motivation: To validate the sustained performance of Headless-AD, we progressed to a more complex Multi-Armed Bandit (MAB) extension that integrated states and real-valued rewards.
Setup: Each time step presented a context with arm features modeled as two-dimensional vectors, and rewards were generated with a standard deviation $\sigma=1$. The data were created using a LinUCB (Li et al., 2010) algorithm trained over $300$ steps, on bandits with $4-20$ arms. We used $100$ contextual bandits, $5$ seeds and $300$ evaluation steps.
Results and Discussion: As shown in Figure 6, Headless-AD’s performance is on par with LinUCB across varied arm counts. While AD also reaches the performance of LinUCB on lower space sizes, it has problems converging on larger action space sizes, in addition to requiring a specific retraining. This serves to highlight the advantages of Headless-AD for use in even more complex environments than toy Bernoulli bandits.
@@ 4.3 Darkroom
Setup: The Darkroom environment, inspired by Chevalier-Boisvert et al. (2018) and Jain et al. (2020), consists of a $N\times N$ grid where the agent needs to reach a specific cell for a reward. In our experiment, the action space consisted of $3$-step sequences of $5$ atomic actions: up, down, left, right, noop. As a result, the environment offered $5^{3}$ possible actions. The agent earned a reward of $1$ if the trajectory induced by the action sequence passed through a goal cell, after which the episode finished. Otherwise, the reward was 00. As an observation, the environment offered only the current coordinates of the agent, so the goal information could only be obtained from the agent’s memory. We divided the goals into disjoint sets used for training and testing in order to evaluate Headless-AD’s in-context learning abilities. Furthermore, the action set was randomly split into train and test sets, each including $50$ and $75$ actions respectively, to create five distinct spaces for assessing various generalization aspects. These action sets are visualized in Figure 1.
@@
Train Actions. Comprising the training split, this set represented the actions encountered during model training.
@@
Test Actions. Comprising the test split, this set assessed Headless-AD’s generalization on novel and larger action sets.
@@
All Actions. Combining both training and testing actions, this set contained $125$ actions and was $2.5$ times larger than the training set. Its aim was to challenge Headless-AD to effectively integrate seen and unseen actions.
@@
Permuted Train Actions. Shuffled training set, meant to test the model’s adaptability to reordered action spaces that comprised of the same actions.
@@
Sliced Test Actions. Tailored to match the training set size, this set contained a slice of the first $50$ actions from the test set. Its aim was to check the models’ generalization abilities on unseen actions while maintaining the action set size.
@@
In scenarios where the action space exceeded the size of the training set, we analyzed the performance of an AD model retrained from scratch. The data generation algorithm was Q-learning, executed over $200$ episodes for each environment. Further details on model hyperparameters are available in Appendix J.
@@
Results and Discussion: Figure 7 illustrates the generalization abilities of both AD and Headless-AD. First, note that both models maintain their performance when transitioning to novel tasks, thus fulfilling their purpose as ICL-RL models. However, this environment was mainly designed to challenge the models’ generalization abilities on novel action spaces. Following the scope of the action space novelty we set earlier, we studied the models’ ability to address changed action semantics and variable action set sizes.
@@
While AD achieves high performance on the train action set, its limitations become evident when the action space is changed. The first limitation is AD’s action set size constraint. The linear layer at the end of its network fixes the size of the output dimension, something that cannot be modified once the model is trained. Increasing the amount of options, as was done in the Test Actions and All Actions sets, requires reinitializing the output dimension and thus retraining the model, which demands additional time and resources. In contrast, once trained, Headless-AD easily adapts to changes in the action set size without losing performance.
Table 1: Ablations: Table compares the performance of Headless-AD with its ablated versions. Columns ’Bandit’ and ’Darkroom’ show the performance averaged along all action sets. Bernoulli Bandit performance is normalized, where 00 denotes a random agent and $1$ denotes the Thompson Sampling algorithm. Columns ’Bandit. Arms Used’ and ’Darkroom. Arms Used’ show the amount of actions tried by the model during evaluation, also averaged along all the action sets. Columns ’Bandit. N Arms’ show the performance of models on the Bernoulli Bandit environment for each respective number of arms during evaluation without averaging. The results are aggregated over $5$ random seeds. As the table shows, changing each component of Headless-AD’s architecture greatly damages the model’s ability either to utilize the action set effectively or to perform well.
Setting Bandit Bandit. Arms Used Darkroom Darkroom. Acts Used Bandit. 20 Arms Bandit. 30 Arms Bandit. 40 Arms Bandit. 50 Arms
Headless-AD 0.98 ± 0.02 19.27 ± 9.26 0.87 ± 0.03 57.21 ± 21.89 0.97 ± 0.02 0.98 ± 0.02 1.0 ± 0.02 1.02 ± 0.04
Prompt Ablation 0.81 ± 0.12 22.28 ± 13.11 0.83 ± 0.03 54.69 ± 18.63 0.91 ± 0.03 0.81 ± 0.05 0.66 ± 0.03 0.55 ± 0.03
Loss Ablation 0.63 ± 0.08 2.4 ± 0.22 0.25 ± 0.04 70.0 ± 29.15 0.6 ± 0.22 0.64 ± 0.27 0.7 ± 0.22 0.8 ± 0.25
Embed Ablation 0.79 ± 0.14 23.78 ± 12.54 0.81 ± 0.03 55.72 ± 19.29 0.85 ± 0.01 0.74 ± 0.03 0.62 ± 0.04 0.52 ± 0.04
@@
The second limitation is AD’s reliance on a stationary action space structure. Due to the classifier nature of AD’s network, it learns to associate each output dimension with a specific action meaning. When action semantics change either due to a permutation of seen actions, as in Permuted Train Actions, or due to a substitution of completely new actions, as in Sliced Test Actions, AD’s performance degrades. We checked whether this problem may be solved by biasing the training distribution to have a structure resembling the one during testing. We trained AD on permuted action sets while leaving the data and the model architecture the same. However, as one can see in Figure 3, this training procedure resulted only in a slightly decreased performance of AD in the Permuted Train setting. An additional graph depicting the performance on each of the (action set, goal type) pairs can be found in Appendix C. Meanwhile, Permuted Train Actions does not pose a challenge for Headless-AD as, by design, it is invariant to specific action order. Most importantly, Headless-AD maintains its performance even when completely new actions are introduced, despite having never seen them in training. We attribute this feature to our success in making the model infer action meaning from the context.